OpenStudy (anonymous):

In a number of two digits, the sum of the digits is 5. If the digits change their places, the new number so formed increases in value by 9. Find the number. Help me solve this problem please.

OpenStudy (anonymous):

2 and 3

OpenStudy (anonymous):

Let us call the digits O and T (for ones and tens). O + T = 5 O*10+T = T*10+O+9 Solve for O and T.

OpenStudy (mr.math):

Let the number be xy, then: \(x+y=5\) \(10x+y=10y+x-9\).

OpenStudy (mr.math):

Solve the system for x and y.

OpenStudy (mr.math):

x=2, y=2. So the number is 23.

OpenStudy (anonymous):

Whilst all that's correct... this is at a pretty basic level ;) I think simultaneous equations are a bit of overkill. 5 doesn't have many factors, assuming we're talking in integers. 23: 2+3 = 5. 32: 23+9.

OpenStudy (anonymous):

Well, but showing how to setup the problem solution using equations is the real purpose of the question. Not necessarily how trivial the answer result is.

OpenStudy (anonymous):

True. And a most excellent answer it was.

OpenStudy (anonymous):

Thank you.